Minden Company is a wholesale distributor of premium European chocolates. The companys balance sheet on April 30 is as folloRequired: 1. Prepare a cash budget for May. (Any Repayments and Interest should be indicated by a minus sign.) MINDEN COM2. Prepare a budgeted income statement for May. MINDEN COMPANY Budgeted Income Statement For the Month Ended May 31 Cost of g3. Prepare a budgeted balance sheet as of May 31. MINDEN COMPANY Budgeted Balance Sheet as of May 31 Assets Total assets Liab
Answer #1

1.

Cash Budget (May)
Beginning Cash Balance $           15,400
Collections from Customers $         400,000 =62000+156000+(520000-156000)*1/2
Total Cash available $         415,400
Deduct : Disbursements
Cash paid to Suppliers $         124,800 =312000*40%
Accounts Payable $           72,600
Operating Expenses $         187,200
New refrigerating equipment $             9,700
Total Cash Disbursements $         394,300
Excess of Receipts over disbursements $           21,100
Financing
Borrowing $           52,000
Repayment $         -17,700
Interest paid $               -100
Total Financing $           34,200
Ending Balance $           55,300

2.

Budgeted Income Statement (May)
Sales Revenue $            520,000
Cost of Goods Sold
   Beginning Inventory $           34,800
   Purchases $         312,000
   Cost of Goods Sold $         346,800
   Less Ending Inventory $         104,000
Cost of Goods Sold $            242,800
Gross profit $            277,200
Operating expenses $            192,400 =187200+5200
Operating Income $               84,800
Interest $                     100
Net Income $               84,700

3.

Budgeted Balance Sheet (May)
Assets
Cash $           55,300
Accounts Receivables $         182,000 =+(520000-156000)*1/2
Inventory $         104,000
Building and Equipment, net of depreciation $         243,500 =239000+9700-5200
Total Assets $         584,800
Liabilities and Stockholder's Equity
Accounts Payable $         187,200 =312000*60%
Notes Payable $           52,000
Capital Shares, no par $         212,000
Retained Earnings $         133,600 =48900+84700
Total Liabilities and Stockholder's Equity $         584,800